# Probabilistically Tightened Linear Relaxation-based Perturbation Analysis for Neural Network Verification

### v1

@@ 3 Probabilistically Tightened LiRPA via Underestimation
Moreover, for vectors ${{\bm{A}}}^{\prime}={{\bm{A}}}^{T}D*$ and ${\bm{b}}^{\prime}={\bm{b}}+{\bm{A}}^{T}{\bm{b}}^{*}$ with probability $\geq\displaystyle{\psi^{\frac{d_{i}+d_{i+1}+\cdots+d_{N}}{m}},}$ it holds that for any $n\in\mathbb{N}$ in any sequence $X$ of $n$ points uniformly and independently sampled from $\mathcal{C},$ for at least $n\times\left(1-\frac{1-R}{m}\sum_{j=i}^{N}d_{j}\right)$ points ${\bm{x}}$ in $X$ we have that
@@
The proof closely follows the analogous result at the basis of the LiRPA approach Zhang et al. [54]. The only (crucial) difference is that we construct $\underline{{\bm{D}}}^{*}$ and $\underline{{\bm{b}}}^{*}$, i.e., the diagonal matrix, and the bias vector meant to provide a linear lower bound on the post-activation values of layer $i,$ using the vectors of estimated reachable sets of the nodes in the layer, instead of their actual reachable sets.
@@
Because of the way we compute the estimated reachable sets $[\overline{l}_{j},\underline{u}_{j}]$ (Proposition 3.2) we have that for each $j$
@@
Moreover these events are independent, hence
@@
In words, we have shown that with probability $\psi^{1/m}$ for each point ${\bm{x}}$ in a fraction of $X$ of size $(1-\frac{(1-R)d_{i}}{m})|X|,$ for all nodes $z^{(i)}_{j}$ in layer $i,$ the pre-activation value $z^{(i)}_{j}({\bm{x}})$ induced by ${\bm{x}}$ is contained in the estimated reachable sets $[\overline{l}_{j},\underline{u}_{j}].$
In order to simplify the notation, let us fix an ${\bm{x}}$ from $X_{OK}$ and let ${\bm{v}}={\bm{z}}^{(i)}({\bm{x}})$ be the vector of pre-activation values induce by ${\bm{x}}$ in the nodes of layer $i.$
We will show that for the $\underline{{\bm{D}}}^{*}$ and $\underline{{\bm{b}}}^{*}$ defined in the statement, we have that for each $j$ it holds that ${{\bm{A}}_{j}}(ReLU({\bm{v}}_{j}))\geq{{\bm{A}}_{j}}(\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j}),$ which immediately implies the desired result, since ${{\bm{A}}}^{T}ReLU({\bm{v}})=\sum_{j}{{\bm{A}}_{j}}(ReLU({\bm{v}}_{j}))\geq{{\bm{A}}_{j}}(\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j})={{\bm{A}}}^{T}(\underline{{\bm{D}}}^{*}{\bm{v}}_{{\bm{x}}}+\underline{{\bm{b}}}^{*}).$
@@
where $[{\bm{l}}^{*}_{j},{\bm{u}}^{*}_{j}]$ is the actual reachable set for the $j$-th node of the layer under consideration.
@@
From inequality 6, we know that since we are underestimating true bounds $[{\bm{l}}^{*}_{j},{\bm{u}}^{*}_{j}]$, the ReLU node is actually unstable, even for any LiRPA approach. Comparing the diagonal coefficients $\frac{\underline{{\bm{u}}}_{j}}{\underline{{\bm{u}}}_{j}-\overline{{\bm{l}}}_{j}}$ with $\frac{{\bm{u}}_{j}}{{\bm{u}}_{j}-{\bm{l}}_{j}}$ and the biases $-\frac{\underline{{\bm{u}}}_{j}\overline{{\bm{l}}}_{j}}{\underline{{\bm{u}}}_{j}-\overline{{\bm{l}}}_{j}}$ with $-\frac{{\bm{u}}_{j}{\bm{l}}_{j}}{{\bm{u}}_{j}-{\bm{l}}_{j}}$ of PT-LiRPA and any LiRPA cannot be helpful. The relation between the coefficients strongly depends on the quality of the bounds computed, and we cannot draw any direct conclusion since in some cases $\underline{{\bm{D}}}>\underline{{\bm{D}}}^{*}$ and in some cases not. Hence, we need to proceed by subcases.
@@
where the last inequality holds since ${\bm{v}}_{j}\in[\overline{{\bm{l}}}_{j},\underline{u}_{j}]$ implies ${\bm{v}}_{j}-\overline{{\bm{l}}}_{j}\geq 0.$
Multiplying both sides by ${\bm{A}}_{j}<0$ and using $ReLU({\bm{v}}_{j})=0,$ we get
@@
where the last inequality holds since under the standing hypothesis the coefficient of ${\bm{v}}_{j}$ in the left hand side is $\geq 1$ and term $-\frac{\underline{{\bm{u}}}_{j}\overline{{\bm{l}}}_{j}}{\underline{{\bm{u}}}_{j}-\overline{{\bm{l}}}_{j}}$ is non-negative.
Multiplying both sides by ${\bm{A}}_{j}<0$ and using $ReLU({\bm{v}}_{j})={\bm{v}}_{j},$ we get
@@
$ReLU({\bm{v}}_{j})=0\geq\alpha_{j}{\bm{v}}_{j}+0=\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j}$
, since $0<\alpha_{j}.$ Hence, multiplying both sides by ${\bm{A}}_{j}\geq 0$ we obtain again the desired inequality.
Subcase 1.4. ${\bm{A}}_{j}\geq 0$ and ${\bm{v}}_{j}\geq 0$.
@@
since $\alpha<1.$ Again, multiplying both sides by ${\bm{A}}_{j}\geq 0$ we obtain again the desired inequality.
This concludes the first case.
$z^{(i)}_{j}$ outside the reachable set $[\overline{l}_{j},\underline{u}_{j}]$. Thus, with probability $\geq\psi$.
@@
Since ${\bm{v}}_{j}\in[\overline{l}_{j},\underline{u}_{j}]$ and $\underline{u}_{j}<0$ we have $ReLU({\bm{v}}_{j})=0.$ Moreover, under the standing hypothesis, we also have $\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j}=0=ReLU({\bm{v}}_{j}).$ Again, multiplying both sides by ${\bm{A}}_{j}$ we have that the desired inequality holds also in this case.
We have proved that for each ${\bm{x}}$ in $X_{OK}$ we have ${\bm{A}}^{T}ReLU({\bm{v}}_{j})\geq{\bm{A}}^{T}(\underline{{\bm{D}}}^{*}{\bm{v}}_{x}+\underline{{\bm{b}}}^{*}).$ Finally, recalling that with probability $\psi^{d_{i}/m}$ we have that $|X_{OK}|\geq|X|(1-\frac{1-R}{m}d_{i}),$ we have that the previous inequality holds with the desired statistical guarantees, which completes the proof of the first statement.
For the second statement it is enough to note that the guarantees on ${\bm{A}},{\bm{b}}$ and (3) hold independently of the choice of the samples leading to the statistical guarantee on (4). Hence they simultaneously happen–yielding that (5) holds–with the product of the probabilities of (3) and (4), i.e., ${\psi^{\frac{d_{i}+d_{i+1}+\cdots+d_{N}}{m}}}.$ In particular, with this probability, for any $n\in\mathbb{N}$ in any sequence $X$ of $n$ points uniformly and independently sampled from $\mathcal{C},$ neither of (3) and (4) fails on at least $n\times\left(1-\frac{1-R}{m}\sum_{j=i}^{N}d_{j}\right)$ points.
@@
The proof then directly follows from Lemma 3.3 and the derivations of Zhang et al. [54]. ∎
@@ 3.1 A qualitative bound of statistical prediction of tolerance limit
The result of [50] allows us to bound the error of the sample-based procedure employed in our PT-LiRPA framework in terms of the number of potential violations in future samples. However, it does not provide any information on the magnitude of such possible violations with respect to the estimated bounds. Additionally, the probabilistic guarantees we provide so far are, broadly speaking, weaker than those offered by related probabilistic methods such as [48, 12], where the guarantees hold for any input ${\bm{x}}\in\mathcal{C}$, and not only for a predefined fraction of it. In this section, we aim to complement Wilks [50] statistical result with a qualitative interpretation that bounds the potential error between the true minimum and its estimate based on samples. Specifically, we leverage known results on extreme value theory [19] to strengthen and extend our original guarantee to apply to the entire perturbation set $\mathcal{C}$, thus bringing the proposed solution in line with other state-of-the-art probabilistic approaches.
We start by noticing that the problem of finding an empirical minimizer close to a function’s exact minimizer is well established in the statistical literature (see, e.g., [2]). Exploiting quantitative assumptions on the objective function—such as a Lipschitz condition that holds in our setting—can facilitate the bounding of the possible magnitude error in the sampled-based minimizer computation approach. In this vein, we begin by providing a first worst-case qualitative bound on the maximum $\Delta$ error we can achieve when employing our sampling-based approach to compute the estimated reachable set with respect to the real (unknown) ones. For readability, in the following we simplify the notation $z_{j}^{(i)}$—which denotes the pre-activation value of the node in position $j$ of layer $i$—by removing the indices and referring generically to a node’s pre-activation value as $z$. We adopt the same simplification for the corresponding lower and upper bound estimates, denoted as $\overline{l}$ and $\underline{u}$, respectively.
@@
$\leq Pr[\forall k\;L||{\bm{x}}_{k}-{\bm{x}}^{*}||_{2}\geq\Delta]$
$=Pr[\forall k\;||{\bm{x}}_{k}-{\bm{x}}^{*}||_{2}\geq\frac{\Delta}{L}]$
Fix $S=\{{\bm{x}}\in\mathcal{C}:||{\bm{x}}-{\bm{x}}^{*}||_{2}\leq\Delta/L\}.$ This is the sphere of radius $\Delta/L$ centered at ${\bm{x}}^{*}$. For a uniformly sampled point ${\bm{x}}\in\mathcal{C},$ let $\mu(S)=Pr[{\bm{x}}inS].$ This is equal to the volume of the set $S$ which is proportional to $(\Delta/L)^{d}\cdot c(d)$, where $c(d)=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is a constant that depends on the dimension $d$. Hence, by the independence of the ${\bm{x}}_{k}$s, the probability $Pr[\forall k\;{\bm{x}}_{k}\notin S]=(1-\mu(S))^{n}$.
$Pr[|\overline{l}-l^{*}|\geq\Delta]$ $\leq Pr[\forall k\;||{\bm{x}}_{k}-{\bm{x}}^{*}||_{2}\geq\frac{\Delta}{L}]$
@@
Clearly, the bound given in Theorem 3.5 accounts for the worst-case scenario and assumes no specific property of the distribution on the input. When the perturbation region has high dimensionality or the DNN has a large Lipschitz constant, this bound may become too loose to offer meaningful insights. Following EVT, and particularly the results of [15], we can provide a tighter bound on the error in the estimation.
@@
Importantly, Lemma 3.6, holds under the assumption that the samples $Y_{1},\ldots,Y_{n}$ follow a nondegenerate limit distribution function in the form $1-exp(-x^{a})$ for some $a>0$ [15]. In particular, this is true for uniformly differentiable $f$, which is the case for most common neural networks, e.g., when the activation functions are Sigmoid, Tanh, etc. For ReLU-based networks, where the differentiability requirement is not satisfied, we can still apply Lemma 3.6 by splitting the non-linearities, i.e., splitting all the ReLU nodes, achieving a linear system differentiable everywhere. Specifically, when a ReLU split introduces new linear constraints, we apply rejection sampling to ensure that drawn samples satisfy the additional constraints. Since the points sampled from the original perturbation region are i.i.d., the rejection phase acts simply as a filter to enforce the constraints, still resulting in a computationally efficient method.

### v2

@@ 3 Probabilistically Tightened LiRPA via Underestimation
  2. 2.
    with probability $\geq\displaystyle{\psi^{\frac{d_{i}+d_{i+1}+\cdots+d_{N}}{m}}}$, for vectors ${{\bm{A}}}^{\prime}={{\bm{A}}}^{T}{\bm{D}}^{*}$ and ${\bm{b}}^{\prime}={\bm{b}}+{\bm{A}}^{T}{\bm{b}}^{*}$ it holds that for any $n\in\mathbb{N}$ in any sequence $X$ of $n$ points uniformly and independently sampled from $\mathcal{C},$ for at least $n\times\left(1-\frac{1-R}{m}\sum_{j=i}^{N}d_{j}\right)$ points $\bm{x}$} in $X$ we have that
@@
The proof closely follows the analogous result at the basis of the LiRPA approach of crown. The only (crucial) difference is that we construct $\underline{{\bm{D}}}^{*}$ and $\underline{{\bm{b}}}^{*}$, i.e., the diagonal matrix, and the bias vector meant to provide a linear lower bound on the post-activation values of layer $i,$ using the vectors of estimated reachable sets of the nodes in the layer, instead of their actual reachable sets.
@@
Because of the way we compute the estimated reachable sets $\overline{l}_{j},\underline{u}_{j}]$ (Proposition 3.2) we have that for each $j$
@@
Moreover, these events are independent, hence
@@
In words, we have shown that with probability $\psi^{1/m}$ for each point $\bm{x}$} in a fraction of $X$ of size $(1-\frac{(1-R)d_{i}}{m})|X|,$ for all nodes $z^{(i)}_{j}$ in layer $i,$ the pre-activation value $z^{(i)}_{j}({\bm{x}})$ induced by $\bm{x}$} is contained in the estimated reachable sets $\overline{l}_{j},\underline{u}_{j}].$
In order to simplify the notation, let us fix an $\bm{x}$} from $X_{OK}$ and let ${\bm{v}}={\bm{z}}^{(i)}({\bm{x}})$ be the vector of pre-activation values induced by $\bm{x}$} in the nodes of layer $i.$ We will show that for the $\underline{{\bm{D}}}^{*}$ and $\underline{{\bm{b}}}^{*}$ defined in the statement, we have that for each $j$ it holds that ${{\bm{A}}_{j}}(ReLU({\bm{v}}_{j}))\geq{{\bm{A}}_{j}}(\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j}),$ which immediately implies the desired result, since ${{\bm{A}}}^{T}ReLU({\bm{v}})=\sum_{j}{{\bm{A}}_{j}}(ReLU({\bm{v}}_{j}))\geq{{\bm{A}}_{j}}(\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j})={{\bm{A}}}^{T}(\underline{{\bm{D}}}^{*}{\bm{v}}_{{\bm{x}}}+\underline{{\bm{b}}}^{*}).$
@@
where ${\bm{l}}^{*}_{j},{\bm{u}}^{*}_{j}]$ is the actual reachable set for the $j$-th node of the layer under consideration.
@@
From inequality 7, we know that since we are underestimating true bounds ${\bm{l}}^{*}_{j},{\bm{u}}^{*}_{j}]$, the ReLU node is actually unstable, even for any LiRPA approach. Comparing the diagonal coefficients $\frac{\underline{{\bm{u}}}_{j}}{\underline{{\bm{u}}}_{j}-\overline{{\bm{l}}}_{j}}$ with $\frac{{\bm{u}}_{j}}{{\bm{u}}_{j}-{\bm{l}}_{j}}$ and the biases $\frac{\underline{{\bm{u}}}_{j}\overline{{\bm{l}}}_{j}}{\underline{{\bm{u}}}_{j}-\overline{{\bm{l}}}_{j}}$ with $\frac{{\bm{u}}_{j}{\bm{l}}_{j}}{{\bm{u}}_{j}-{\bm{l}}_{j}}$ of PT-LiRPA and any LiRPA cannot be helpful. The relation between the coefficients strongly depends on the quality of the bounds computed, and we cannot draw any direct conclusion since in some cases $\underline{{\bm{D}}}>\underline{{\bm{D}}}^{*}$ and in some cases not. Hence, we need to proceed by subcases.
@@
where the last inequality holds since ${\bm{v}}_{j}\in[\overline{{\bm{l}}}_{j},\underline{u}_{j}]$ implies ${\bm{v}}_{j}-\overline{{\bm{l}}}_{j}\geq 0.$ Multiplying both sides by ${\bm{A}}_{j}<0$ and using $ReLU({\bm{v}}_{j})=0,$ we get
@@
where the last inequality holds since under the standing hypothesis the coefficient of ${\bm{v}}_{j}$ in the left hand side is $\geq 1$ and term $\frac{\underline{{\bm{u}}}_{j}\overline{{\bm{l}}}_{j}}{\underline{{\bm{u}}}_{j}-\overline{{\bm{l}}}_{j}}$ is non-negative. Multiplying both sides by ${\bm{A}}_{j}<0$ and using $ReLU({\bm{v}}_{j})={\bm{v}}_{j},$ we get
@@
$ReLU({\bm{v}}_{j})=0\geq\alpha_{j}{\bm{v}}_{j}+0=\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j}\;,$
since $0<\alpha_{j}.$ Hence, multiplying both sides by ${\bm{A}}_{j}\geq 0$ we obtain again the desired inequality. Subcase 1.4. ${\bm{A}}_{j}\geq 0$ and ${\bm{v}}_{j}\geq 0$.
@@
since $\alpha<1.$ Again, multiplying both sides by ${\bm{A}}_{j}\geq 0$ we obtain again the desired inequality. This concludes the first case.
@@
Since ${\bm{v}}_{j}\in[\overline{l}_{j},\underline{u}_{j}]$ and $\underline{u}_{j}<0$ we have $ReLU({\bm{v}}_{j})=0.$ Moreover, under the standing hypothesis, we also have $\underline{{\bm{D}}}^{*}_{j,j}{\bm{v}}_{j}+\underline{{\bm{b}}}^{*}_{j}=0=ReLU({\bm{v}}_{j}).$ Again, multiplying both sides by ${\bm{A}}_{j}$, we have that the desired inequality holds also in this case. We have proved that for each $\bm{x}$} in $X_{OK}$ we have ${\bm{A}}^{T}ReLU({\bm{v}}_{j})\geq{\bm{A}}^{T}(\underline{{\bm{D}}}^{*}{\bm{v}}_{x}+\underline{{\bm{b}}}^{*}).$ Finally, recalling that with probability $\psi^{d_{i}/m}$ we have that $|X_{OK}|\geq|X|(1-\frac{1-R}{m}d_{i}),$ we have that the previous inequality holds with the desired statistical guarantees, which completes the proof of claim (1).
For proving claim (2) it is enough to note that the guarantees on ${\bm{A}},{\bm{b}}$ and (4) hold independently of the choice of the samples leading to the statistical guarantee on (5). Hence they simultaneously happen–yielding that (6) holds–with the product of the probabilities of (4) and (5), i.e., ${\psi^{\frac{d_{i}+d_{i+1}+\cdots+d_{N}}{m}}}.$ In particular, with this probability, for any $n\in\mathbb{N}$ in any sequence $X$ of $n$ points uniformly and independently sampled from $\mathcal{C},$ neither of (4) and (5) fails on at least $n\times\left(1-\frac{1-R}{m}\sum_{j=i}^{N}d_{j}\right)$ points.
@@
The proof then directly follows from Lemma 3.3 and the derivations of crown. ∎
@@ 3.1 A Qualitative Bound of Statistical Prediction of Tolerance Limit
The result of (wilks) allows us to bound the error of the sample-based procedure employed in our PT-LiRPA framework in terms of the number of potential violations in future samples. However, it does not provide any information on the magnitude of such possible violations with respect to the estimated bounds. Additionally, the probabilistic guarantees we provide so far are, broadly speaking, weaker than those offered by related probabilistic methods such as (proven; randomizedSmoothing), where the guarantees hold for any input ${\bm{x}}\in\mathcal{C}$, and not only for a predefined fraction of it. In this section, we aim to complement wilks statistical result with a qualitative interpretation that bounds the potential error between the true minimum and its estimate based on samples. Specifically, we leverage known results on extreme value theory (haan2006extreme) to strengthen and extend our original guarantee to apply to the entire perturbation set $\mathcal{C}$, thus bringing the proposed solution in line with other state-of-the-art probabilistic approaches.
We start by noticing that the problem of finding an empirical minimizer close to a function’s exact minimizer is well established in the statistical literature (see, e.g., (archetti1984survey)). Exploiting quantitative assumptions on the objective function—such as a Lipschitz condition that holds in our setting—can facilitate the bounding of the possible magnitude error in the sampled-based minimizer computation approach. In this vein, we begin by providing a first worst-case qualitative bound on the maximum $\Delta$ error we can achieve when employing our sampling-based approach to compute the estimated reachable set with respect to the real (unknown) ones. For readability, in the following we simplify the notation $z_{j}^{(i)}$—which denotes the pre-activation value of the node in position $j$ of layer ii—by removing the indices and referring generically to a node’s pre-activation value as $z$. We adopt the same simplification for the corresponding lower and upper bound estimates, denoted as $\overline{l}$ and $\underline{u}$, respectively.
@@
[…]
@@
Figure 4: Illustrative representation in 3D of Theorem 3.5.
Broadly speaking, in the theorem, we bound the probability that the estimated minimizer deviates from the true minimizer by more than a predefined threshold $\Delta$. Owing to the Lipschitz properties of DNNs, this is equivalent to estimating the probability that a random sample of inputs includes at least one point within distance $\Delta/L$ of the true minimizer $\mathbf{x}^{*}$, i.e., an input $\bm{x}$} in $S=\{\mathbf{x}\in\mathcal{C}:\|\mathbf{x}-\mathbf{x}^{*}\|_{2}\leq\Delta/L\}$. Geometrically, $S$ corresponds to the intersection of the perturbation region $\mathcal{C}$ with a ball of radius $\Delta/L$ centered at $\mathbf{x}^{*}$. If $\mathbf{x}^{*}$ lies sufficiently far from the boundary of $\mathcal{C}$, then $S$ is exactly such a ball. Otherwise, if $\mathbf{x}^{*}$ is near the boundary, $S$ becomes a truncated ball, i.e., the intersection with $\mathcal{C}$. This is clearly represented in Fig. 4, where we depict a potential perturbation region $\mathcal{C}$ in 3d derived from the original input ${\bm{x}}_{0}$ and the minimizer $\bm{x}$$^{*}$ of $f$ for $\mathcal{C}$.
Clearly, the bound given in Theorem 3.5 accounts for the worst-case scenario and assumes no specific property of the distribution on the input. When the perturbation region has high dimensionality or the DNN has a large Lipschitz constant, this bound may become too loose to offer meaningful insights. Following EVT, and particularly the results of (de1981estimation), we can provide a tighter bound on the error in the estimation.
@@
Importantly, Lemma 3.6, holds under the assumption that the samples $Y_{1},\ldots,Y_{n}$ follow a nondegenerate limit distribution function in the form $1-exp(-x^{a})$ for some $a>0$ (de1981estimation). In particular, this is true for uniformly differentiable $f$, which is the case for most common neural networks, e.g., when the activation functions are Sigmoid, Tanh, etc. For ReLU-based networks, where the differentiability requirement is not satisfied, we can still apply Lemma 3.6 by splitting the non-linearities, i.e., splitting all the ReLU nodes, achieving a linear system differentiable everywhere babSplitRelu.
Concretely, whenever a ReLU’s interval bound crosses 0 (i.e., $\underline{z}<0<\overline{z}$) computed over the perturbation region $\mathcal{C}$, we split the computation into two linear branches enforcing $z({\bm{x}})\geq 0$ (active) and $z({\bm{x}})\leq 0$ (inactive). On each resulting subregion, the network is affine, so the neuron’s pre-activation is an affine (hence differentiable) map of $\bm{x}$}. Sampling i.i.d. uniformly from $\mathcal{C}$ and rejecting points that do not satisfy a branch’s linear constraints is equivalent to sampling i.i.d. uniformly from that constrained subregion. Hence, the order statistics restricted to the subregion obey the required extreme-value limit. In practice, in our experimental setting described in Sec. 5, where $\mathcal{C}$ is an $\ell_{\infty}$-ball and ReLU constraint splitting is applied, the rejection rate remains very low. Specifically, the fraction of discarded samples is always negligible (typically below 5% on average). This is because the probability that a uniformly sampled point violates the constraints decreases rapidly as the sample size increases, and the rejection step only eliminates points outside the feasible region while preserving the independence of the remaining samples.